The armature of an ac generator is a rectangular coil 2.00 cm by 6.00 cm with 80 turns. It is immersed in a uniform magnetic field of magnitude 0.250 T. If the amplitude of the emf in the coil is 22.0 V, at what angular speed is the armature rotating?
Answer #1

flux through the coil is given by

\phi = NBA cos(\theta)

induced emf can be calculated as

\epsilon =- \frac{d\phi}{dt}

\epsilon =NBA sin(\theta)\frac{d\theta}{dt}

since \frac{d\theta}{dt} =\omega

So,

\epsilon =NBA \omega sin(\theta)

maximum emf when \theta = 90

\epsilon =NBA \omega

22 =(80)(0.25)(0.02\times 0.06) \omega

\omega = 1833.33 rad/s